A rectangle has a length os 12cm and a width of 4cm.
a. Find the perimeter.
b. Find the area.
Thank you

Answer by mukhi1(5)

a. Perimeter = 2 * Width + 2* length
= 4*2 + 12*2 = 32cm
b. Area = length * width
= 4 *12
= 48 cm^2